OpenStudy (anonymous):

What is it called, if I double a number and double that result again and double that result again and so on? for example: 3 - 6 - 12 -24 or 7.85 - 15.7 - 31.4 - 62.8 this kind of operation or row must have a specific name, hasn't it? it's not exponentiation and it's also not Multiples of a number. it's....

OpenStudy (shadowlegendx):

I think sequences maybe...

OpenStudy (anonymous):

hmm yeah, sequence could be true, since it could be defined recursively with \[x_n = x_{n-1} * 2\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes this is geometric progession

OpenStudy (anonymous):

\(x_{n+1}=2x_n\implies x_n=2^nx_0=3\cdot2^n\)